OpenStudy (anonymous):

I need help with this I followed the steps presented to me and still got it wrong. Also need full explaination!!!! Here is the question: What is the molar concentration in a solution prepared by dissolving 22g of calcium chloride(FW=111.1g/mol) to 4.0L with water? 22/111.1=.19801 M1V1-M2V2 (.19801)(22)=(M)(4000) 4.35622=M(4000) 4.35622/4000= .001089055 Answer not found among possible answers. Possible Answers: 0.10M 0.63M

OpenStudy (hlilly2413):

Ok, don't freak you just used the wrong equation. So, to find Molarity (M) you use the following equation: \[Molarity (M) =\frac{ moles }{ Liter }\]

OpenStudy (hlilly2413):

You were correct in changing your grams of CaCl2 to moles, so nice job there! You will need to use that value for the equation above. Also, the M1V1=V2M2 equation is typically used when you have two different substances and you know the Molarity and Volume for one, but not both for the other. One last thing, your answer is in the list so long as you remember to round. Hope this helps!

OpenStudy (anonymous):

It really doesn't and I have been working these problems all day so yeah I just about had it!!!

OpenStudy (matt101):

As @hlilly2413 said, you were correct to convert the mass of CaCl2 to moles. The equation for calculating molarity of a solution is exactly as written above, M=n/V. The units of molarity are mol/L, so you'll want to make sure your n is some number of moles, and V is volume in liters. So you have M=0.19801 mol / 4 L = 0.050 M and the first answer is the correct one! To reiterate what @hlilly2413 said, the equation you used, M1V1=M2V2, is generally used when you're performing some sort of dilution and are concerned about both the before and after situations (effectively two different solutions). The moles of solute in solution don't change (hence the equality), but the volume, and therefore the concentration, do. Here we're only dealing with one solution, so we can use the basic M=n/V equation to find what we want! Does that make sense?

OpenStudy (anonymous):

That answer is wrong as well. Just finished with my quiz again and 0.050 M is incorrect. So try again.

OpenStudy (matt101):

The answer for the question you provided is definitely 0.050 M. The only thing I can think of is maybe the question is asking for the molar concentration OF CHLORINE ONLY, in which case the answer would be doubled (0.10 M) since there are two Cl atoms for every molecule of CaCl2. Can you check and see that you copied the question correctly?

OpenStudy (anonymous):

That is how it's laid out. I took it down word for word.

OpenStudy (whpalmer4):

http://www.wolframalpha.com/input/?i=22+g+of+calcium+chloride+in+4+l+water If your quiz program says 0.050 M is incorrect, someone entered the wrong scoring data.

OpenStudy (whpalmer4):

Though I do agree that if the problem is asking for the molar concentration of Cl ions, 0.10 M would be a correct answer. But the problem we have before us does not ask for that — how would we know they want the Cl concentration, and not the Ca concentration (which is 0.050 M), without some language to that effect?

OpenStudy (matt101):

^

OpenStudy (anonymous):

Same way we need to know the answer to everything else in chemistry guess till you get it right.

OpenStudy (anonymous):

Thanks anyway but I pretty much had it tonight with this chemistry non-seance. 8th time failing a quiz so yeah I am done with this BS. TY anyway.

OpenStudy (whpalmer4):

Strange, I don't remember guessing being covered in chemistry class. Maybe it is an more advanced topic.

OpenStudy (hlilly2413):

You're welcome for attempting to help you. I'm sorry that the formula didn't work for this problem. To be fair, you DID have .050 M (or something very similar) as a possible answer choice when I originally posted the assistance. Perhaps ask your professor or a tutor to assist you because the way the question is laid out, that IS your answer. Guessing is not the way to figure these problems out. They are typically systematic. Good luck!